# Privacy-Preserving Video Anomaly Detection: A Survey

### v1

@@ Abstract
Video Anomaly Detection (VAD) aims to automatically analyze spatiotemporal patterns in surveillance videos collected from open spaces to detect anomalous events that may cause harm without physical contact. However, vision-based surveillance systems such as closed-circuit television often capture personally identifiable information. The lack of transparency and interpretability in video transmission and usage raises public concerns about privacy and ethics, limiting the real-world application of VAD. Recently, researchers have focused on privacy concerns in VAD by conducting systematic studies from various perspectives including data, features, and systems, making Privacy-Preserving Video Anomaly Detection (P2VAD) a hotspot in the AI community. However, current research in P2VAD is fragmented, and prior reviews have mostly focused on methods using RGB sequences, overlooking privacy leakage and appearance bias considerations. To address this gap, this article systematically reviews the progress of P2VAD for the first time, defining its scope and providing an intuitive taxonomy. We outline the basic assumptions, learning frameworks, and optimization objectives of various approaches, analyzing their strengths, weaknesses, and potential correlations. Additionally, we provide open access to research resources such as benchmark datasets and available code. Finally, we discuss key challenges and future opportunities from the perspectives of AI development and P2VAD deployment, aiming to guide future work in the field.
@@ I Introduction
Video Anomaly Detection (VAD) aims to automatically identify irregular patterns in spatio-temporal surveillance data to detect unexpected anomalous events [1]. Due to its ability to capture real-time environmental information in open spaces without physical contact, VAD has demonstrated promising applications in emerging fields such as smart cities [2], modern industry [3], and healthcare [4]. Applications include traffic accident warning in Intelligent Transportation Systems (ITS) [5], identifying irregularities in industrial production [6], and detecting elderly falls. Compared to Action Recognition (AR) [7, 8, 9], which relies on fine-grained labels for model training, VAD generally follows the open-world assumption that real-world anomalies are rare and unbounded [10]. As a result, collecting all possible anomalous events in diverse and dynamic scenarios for training supervised multi-class classification models is impractical. Moreover, AR methods are ill-suited to handle the significant data imbalance and label noise often encountered in anomaly detection. In contrast, VAD is often formulated as an unsupervised outlier detection task, using easily collected regular events to train the model to characterize the prototypical patterns of normal videos, while treating uncharacterizable samples as anomalies [11]. Thus, Unsupervised VAD (UVAD) avoids the expensive data preparation cost of collecting anomalies and exhibits great scene adaptability, with the theoretical ability to detect any positive samples distinct from negative ones [12, 13, 14].
In recent years, researchers have proposed Weakly-supervised VAD (WsVAD) approaches, which utilize video-level labeled regression models to compute fine-grained frame-level anomaly scores for the temporal localization of anomalous events [15]. Although such approaches [16, 17, 18] are usually limited to identifying predefined anomaly categories in a given scene, they still do not require segment-level or frame-level labeling, but instead use binary labels (0 or 1) to indicate whether a video contains anomalies or not, without the need for second-level categorization. Existing work [17] has shown that weakly-supervised VAD models do not need to consider the data balance between various types of anomalous samples and regular samples when identifying anomalous behaviors that may threaten life safety or cause economic loss in specific scenarios (e.g., criminal events and violent behaviors). These models have proven to be more reliable and practical than weakly-supervised action recognition models that rely on multi-class labels. The latest VAD research aims to detect anomalies directly from raw, unfiltered surveillance data. The so-called Fully-unsupervised VAD (FuVAD) allows models to be automatically trained using data from large-scale video IoT systems and online video platforms.
@@ I-A Related Surveys
Source Main Focus P2 Pathways Discussed Viewpoint
@@ I-B P2VAD Taxonomy
With a thorough review of recent advance in video anomaly detection, this article introduces the first VAD taxonomy from the privacy-preserving perspective, illustrated in Fig. 1. The proposed P2VAD taxonomy includes all related works across three key stages: S1: date acquisition; S2: model learning, and S3: system applications, which are accordingly divided into three main categories:
  1. 1.
    P2VAD with Non-Identifiable Elements (NIE): NIE methods aim to avoid the capture or storage of privacy-sensitive RGB sequences during data acquisition and pre-procession. The means of privacy preservation include encrypted video coding, Non-Visible Light Cameras (NVLCs), or using object detection models to mask human-related regions in RGB sequences, ensuring that the data for modeling lacks identifiable information.
  2. 2.
    P2VAD with Desensitized Intermediate Modalities (DIM): Since most real-world anomalies are motion-ralated, intermediate modalities that capture temporal cues, such as skeleton keypoints or optical flow, are used during P2VAD model learning. Such data modalities can not only eliminate the privacy concerns related to appearance but also improve the robustness of P2VAD models against pixel noise in RGB videos.
  3. 3.
    Edge-Cloud Intelligence (ECI) Empowered P2VAD: Beyond the risks of appearance privacy breaches from RGB sequences, security concerns can arise during the data transmission in large-scale edge-cloud systems. Techniques like privacy computing, edge-cloud collaboration, and federated learning have been introduced to VAD to address these concerns at the application stage.
Moreover, NIE methods can be further divided based on the format of the final acquired video: 1) Encrypted Compressed Video-based: Formats like H.264 or H.265 are used to compress RGB video into indistinguishable binary streams, allowing deep learning models to extract features directly from the bitstream. This reduces the need for identifying personal information and lowers data transmission costs. 2) Sensitive Region Masking/Synthetic Video-based: Anomalies often involve abnormal human motion, making color and texture information unnecessary. Pre-trained object detection and instance segmentation models can mask human figures in RGB frames to input into privacy-preserving VAD models. Additionally, datasets like UBnormal simulate abnormal behaviors using avatars, avoiding real identifiable information while training privacy-agnostic VAD models. 3) NVLCs Videos-based: Devices like infrared or depth cameras capture spatio-temporal environmental data that lacks texture but reveals contour variations, aiding in understanding anomalous behaviors.
@@
While NIE methods may increase costs related to data preparation and require specialized hardware, DIM approaches utilize RGB sequences and learn normal behavior from intermediate modalities, such as skeleton data or optical flow. Skeleton-based DIM approaches focus on human behavioral anomalies and are more resistant to pixel noise, providing excellent performance in human-centric VAD tasks. These approaches can be classified into four categories: 1) agent task, 2) clustering, 3) distribution modeling, and 4) autoregression. Optical flow, while useful for capturing motion, can also raise privacy concerns when used in conjunction with RGB sequences. We analyze both hybrid and optical flow-alone approaches to demonstrate their potential for P2VAD and to inspire further research in optical flow modeling.
ECI-enabled methods address privacy concerns related to data leakage in large-scale edge computing systems. By processing sensitive data on edge devices and transmitting only desensitized or encrypted features to the cloud, these systems mitigate the risks of data breaches. End-edge-cloud systems work together to protect against malicious data collection, providing a foundation for privacy-preserving VAD in large-scale applications. Although still in their infancy, initial work in areas such as 1) Privacy Computing-based P2VAD, 2) Edge-Cloud Collaboration VAD, and 3) Federated VAD is emerging.
@@ I-C Contribution Summary
  • This is the first survey, to the best of our knowledge, that focuses on privacy-preserving video anomaly detection. We propose a P2VAD taxonomy comprising three major categories with eight subcategories, systematically summarizing the recent advances in P2VAD across data acquisition, model learning, and system applications, while identifying potential research directions.
@@ II-B1 Encrypted Video Encoding
II-B2 Object Detection
Object detection (OD) techniques are employed to detect and localize target sensitive objects (e.g., humans) within video frames. Commonly used OD algorithms, such as the YOLO family [72] (e.g., YOLOv3 [73] and YOLOv3-spp [74]) and Mask R-CNN [75], can quickly recognize human figures in images and protect privacy by masking these pixels. Combined with instance segmentation [76], P2VAD can more precisely mask human-related pixels in videos, minimizing privacy exposure and avoiding unnecessary spatial noise. By transmitting only frames that exclude identifiable information, P2VAD enables effective VAD while preserving privacy.
@@ III P2VAD with Non-Identifiable Elements
Non-Identifiable Elements (NIE)-based P2VAD methods aim to prevent privacy and security risks by removing or abstracting identifiable information during data acquisition and preprocessing stages. The desensitized data is then used as input to the anomaly detection model. Based on the level of morphological abstraction, we classify existing NIE methods into three categories: encrypted videos, sensitive area masking/synthesis, and NVLCs videos.
In spite of privacy, such methods reduce data volume and prevent extraneous appearance information from interfering with model learning. However, these approaches have drawbacks, such as reliance on additional processing (e.g., compression, object detection) or the need for specialized acquisition equipment, which can significantly increase the cost of model deployment. Below, we analyze the details of each NIE method.

### v2

@@ Abstract
Video Anomaly Detection (VAD) aims to automatically analyze spatiotemporal patterns in surveillance videos collected from open spaces to detect anomalous events that may cause harm, such as fighting, stealing, and car accidents. However, vision-based surveillance systems such as closed-circuit television often capture personally identifiable information. The lack of transparency and interpretability in video transmission and usage raises public concerns about privacy and ethics, limiting the real-world application of VAD. Recently, researchers have focused on privacy concerns in VAD by conducting systematic studies from various perspectives including data, features, and systems, making Privacy-Preserving Video Anomaly Detection (P2VAD) a hotspot in the AI community. However, current research in P2VAD is fragmented, and prior reviews have mostly focused on methods using RGB sequences, overlooking privacy leakage and appearance bias considerations. To address this gap, this article is the first to systematically reviews the progress of P2VAD, defining its scope and providing an intuitive taxonomy. We outline the basic assumptions, learning frameworks, and optimization objectives of various approaches, analyzing their strengths, weaknesses, and potential correlations. Additionally, we provide open access to research resources such as benchmark datasets and available code. Finally, we discuss key challenges and future opportunities from the perspectives of AI development and P2VAD deployment, aiming to guide future work in the field.
@@ I Introduction
Video Anomaly Detection (VAD) aims to automatically identify irregular patterns in spatio-temporal surveillance data to detect unexpected anomalous events [1]. Due to its ability to capture real-time environmental information in open spaces, VAD has demonstrated promising applications in emerging fields such as smart cities [2], modern industry [3], and healthcare [4]. Applications include traffic accident warning in Intelligent Transportation Systems (ITS) [5], identifying irregularities in industrial production [6], and detecting elderly falls. Compared to Action Recognition (AR) [7, 8], which relies on fine-grained labels for model training, VAD generally follows the open-world assumption that real-world anomalies are rare and unbounded [9]. As a result, collecting all possible anomalous events in diverse and dynamic scenarios for training supervised multi-class classification models is impractical. Moreover, AR methods are ill-suited to handle the significant data imbalance and label noise often encountered in anomaly detection. In contrast, VAD is often formulated as an unsupervised outlier detection task, using easily collected regular events to train the model to characterize the prototypical patterns of normal videos, while treating uncharacterizable samples as anomalies [10]. Thus, Unsupervised VAD (UVAD) avoids the expensive data preparation cost of collecting anomalies and exhibits great scene adaptability, with the theoretical ability to detect any positive samples distinct from negative ones [11, 12, 13, 14, 15].
In recent years, researchers have proposed Weakly-supervised VAD (WsVAD) approaches, which utilize video-level labeled regression models to compute fine-grained frame-level anomaly scores for the temporal localization of anomalous events [16, 17]. Although such approaches [18, 19, 20] are usually limited to identifying predefined anomaly categories in a given scene, they still do not require segment-level or frame-level labeling, but instead use binary labels (0 or 1) to indicate whether a video contains anomalies or not, without the need for second-level categorization. Existing work [19] has shown that weakly-supervised VAD models do not need to consider the data balance between various types of anomalous samples and regular samples when identifying anomalous behaviors that may threaten life safety or cause economic loss in specific scenarios (e.g., criminal events and violent behaviors). These models have proven to be more reliable and practical than weakly-supervised action recognition models that rely on multi-class labels. The latest VAD research aims to detect anomalies directly from raw, unfiltered surveillance data. The so-called Fully-unsupervised VAD (FuVAD) allows models to be automatically trained using data from large-scale video IoT systems and online video platforms.
@@ I-A Related Surveys
Venue Main Focus P2 Pathways Discussed Viewpoint
@@ I-B P2VAD Taxonomy
Privacy-preserving video anomaly detection emerges from the recognition that conventional VAD systems, while effective in detecting anomalous behaviors, inherently pose significant privacy risks by processing RGB video sequences that contain identifiable human appearance information, facial features, clothing details, and environmental layouts. These privacy concerns have become increasingly critical as surveillance systems expand across public and private spaces, necessitating the development of VAD approaches that can maintain detection effectiveness while protecting individual privacy and complying with data protection regulations.
@@
To address these challenges systematically, this article introduces the first comprehensive VAD taxonomy from a privacy-preserving perspective, as illustrated in Fig. 1. The taxonomy is constructed around the fundamental understanding that privacy risks in VAD systems manifest at different stages of the data lifecycle, requiring distinct mitigation strategies. Privacy breaches can occur during initial data acquisition when cameras capture identifiable information, during model learning when algorithms process sensitive appearance features, and during system deployment when data is transmitted across distributed computing infrastructures.
The proposed P2VAD taxonomy encompasses all related works across three critical stages corresponding to the data lifecycle: S1: data acquisition; S2: model learning; and S3: system applications. These stages naturally lead to three main categories of privacy-preserving approaches:
  • P2VAD with Non-Identifiable Elements (NIE) represents approaches that fundamentally alter the data acquisition process to avoid capturing or storing privacy-sensitive information from the outset. Non-identifiable elements refer to video data representations that have been processed, encoded, or captured in ways that obscure or eliminate personally identifiable information while preserving motion and behavioral patterns necessary for anomaly detection. These methods operate on the principle that preventing the initial capture of sensitive data is more secure than attempting to protect such data after acquisition. NIE techniques include encrypted video coding that compresses RGB sequences into indistinguishable binary streams, deployment of Non-Visible Light Cameras (NVLCs) that capture thermal or depth information without revealing facial features, and application of object detection models to mask human-related regions in RGB sequences during preprocessing.
  • P2VAD with Desensitized Intermediate Modalities (DIM) focuses on the model learning stage by extracting privacy-neutral intermediate representations from video data that capture essential motion and behavioral information while discarding appearance details. Desensitized intermediate modalities are data representations derived from RGB video sequences that preserve temporal and spatial patterns relevant to anomaly detection while removing identifiable visual characteristics. These modalities leverage the insight that most real-world anomalies are motion-related rather than appearance-related, enabling effective detection through representations such as human skeleton keypoints that capture body pose and movement without revealing facial features, clothing, or other identifying characteristics. Optical flow representations that encode pixel-level motion vectors while discarding color and texture information also fall into this category.
  • Edge-Cloud Intelligence (ECI) Empowered P2VAD addresses privacy risks that emerge during the system application stage, particularly in distributed computing environments where video data or derived features must be transmitted, stored, and processed across multiple devices, edge nodes, and cloud servers. These methods recognize that even when initial data acquisition employs privacy-preserving techniques, additional vulnerabilities can arise from data transmission, collaborative processing, and distributed model training. ECI approaches employ cryptographic techniques such as homomorphic encryption and differential privacy, federated learning protocols that enable collaborative model training without centralizing raw data, and edge-cloud collaboration architectures that process sensitive information locally while leveraging cloud resources for computational scaling.
NIE methods can be further categorized based on the format and characteristics of the acquired video data. Encrypted Compressed Video-based approaches utilize standard video compression formats like H.264 or H.265 to transform RGB sequences into compressed bitstreams that appear as indistinguishable binary data to human observers while preserving sufficient information for machine learning algorithms to extract behavioral patterns. Sensitive Region Masking/Synthetic Video-based methods employ pre-trained computer vision models to identify and obscure human figures or other sensitive objects in video frames, or alternatively generate synthetic datasets using avatars and simulated environments that replicate anomalous behaviors without involving real individuals. NVLCs Video-based approaches deploy specialized camera hardware such as infrared thermal cameras or depth sensors that capture environmental and motion information through non-visible light spectra, naturally avoiding the collection of detailed appearance information.
P2VAD methods with DIM, while utilizing RGB sequences during initial capture, focus on extracting intermediate representations that eliminate privacy-sensitive appearance information during the model learning phase. Skeleton-based DIM methods extract human pose keypoints that represent joint positions and limb orientations, enabling the detection of behavioral anomalies through motion pattern analysis while completely discarding facial features, clothing details, and background information. These approaches demonstrate particular effectiveness in human-centric anomaly detection tasks and exhibit improved robustness against illumination changes, weather conditions, and camera viewpoint variations compared to RGB-based methods. Skeleton-based approaches can be further classified into four methodological categories: agent task-based methods that learn normality through reconstruction or prediction tasks, clustering approaches that group motion patterns and identify outliers, distribution modeling techniques that capture statistical properties of normal behaviors, and autoregressive methods that learn temporal dependencies in motion sequences.
ECI-enabled methods address the growing complexity of modern surveillance systems that increasingly rely on distributed computing architectures spanning edge devices, intermediate processing nodes, and cloud-based analytics platforms. These approaches recognize that privacy protection must extend beyond individual data processing to encompass the entire system ecosystem, including data transmission protocols, distributed storage mechanisms, and collaborative learning frameworks. Privacy Computing-based P2VAD employs cryptographic techniques to enable computation on encrypted data without exposing sensitive information. Edge-cloud collaboration VAD distributes processing tasks strategically between local edge devices that handle sensitive data and cloud resources that perform computationally intensive analytics on privacy-protected features. Federated VAD enables multiple surveillance systems or organizations to collaboratively improve anomaly detection models through distributed learning protocols that share model parameters rather than raw data, ensuring that sensitive video information remains localized while benefiting from collective intelligence.
@@ I-C Contribution Summary
  • To the best of our knowledge, this is the first survey that focuses on privacy-preserving video anomaly detection. We propose a P2VAD taxonomy comprising three major categories with eight subcategories, systematically summarizing the recent advances in P2VAD across data acquisition, model learning, and system applications, while identifying potential research directions.
@@ II-B1 Encrypted Video Encoding
II-B2 Object Detection (OD)
OD techniques are employed to detect and localize target sensitive objects (e.g., humans) within video frames. The application of OD in VAD extends beyond privacy preservation to encompass AD approaches that analyze object-level behaviors and interactions. Object-centric VAD methods have demonstrated significant effectiveness in capturing anomalous patterns by focusing on individual objects or their relationships within scenes. For instance, Ionescu et al. [72] introduce object-centric convolutional auto-encoders that encode both motion and appearance information, formulating abnormal event detection as a one-versus-rest binary classification problem where normality clusters are separated from dummy anomalies. Similarly, Wang et al. [73] propose a self-supervised approach that solves decoupled spatio-temporal jigsaw puzzles, treating VAD as a multi-label fine-grained classification problem that captures discriminative appearance and motion features through spatial and temporal puzzle permutations. However, object-centric methods in conventional VAD research inherently raise privacy concerns when processing RGB sequences, as they typically rely on detailed appearance information to distinguish between normal and abnormal object behaviors. This limitation motivates the development of privacy-preserving object detection strategies in P2VAD systems. Commonly used OD algorithms, such as the YOLO family [74] (e.g., YOLOv3 [75] and YOLOv3-spp [76]) and Mask R-CNN [77], can be repurposed to quickly recognize human figures in images and protect privacy by masking these pixels. Combined with instance segmentation [78], P2VAD can more precisely mask human-related pixels in videos, minimizing privacy exposure and avoiding unnecessary spatial noise. By transmitting only frames that exclude identifiable information, P2VAD enables effective VAD while preserving privacy.
@@ II-B5 Optical Flow
II-C Privacy Preservation Mechanisms in P2VAD
While appearance abstraction and desensitization techniques focus on removing or obscuring identifiable information at the data level, system-level privacy preservation mechanisms provide cryptographic and algorithmic protections during data transmission, storage, and processing phases. These mechanisms are particularly crucial in distributed P2VAD systems where video data or intermediate features must traverse multiple devices, edge nodes, and cloud servers. Understanding these foundational techniques is essential for comprehending the P2VAD frameworks discussed in Sec. V.
II-C1 Homomorphic Encryption
Homomorphic encryption enables computation directly on encrypted data without requiring decryption, which has emerged as a fundamental technique for privacy-preserving anomaly detection in cloud environments. In P2VAD applications, this cryptographic approach allows cloud servers to perform anomaly detection algorithms on encrypted video features or motion vectors while maintaining complete data confidentiality. The mathematical foundation ensures that operations performed on encrypted data yield equivalent results to those performed on plaintext data, albeit in encrypted form. When multiple cameras transmit encrypted motion vector data to a central server for cross-scene anomaly analysis, homomorphic encryption enables the computation of aggregated anomaly scores without exposing raw motion information. Cheng et al. [91] demonstrate this principle through their SecureAD framework, which implements additive secret-sharing protocols to distribute encrypted video features across multiple computing nodes. This approach ensures that no single entity can reconstruct the original sensitive information, effectively addressing privacy concerns in collaborative VAD scenarios.
II-C2 Differential Privacy
Differential privacy provides mathematical guarantees against individual data identification by introducing carefully calibrated noise into the computation process. This technique has proven particularly valuable in P2VAD systems where multiple surveillance cameras contribute data for collaborative anomaly detection. The core mechanism involves adding statistical noise to query results or intermediate computations such that the presence or absence of any single video sequence has negligible impact on the final output. Giorgi et al. [92] apply differential privacy in edge-cloud VAD architectures, where local edge devices transmit noisy feature vectors to cloud aggregators, preventing the reconstruction of information about specific individuals or locations. The noise injection process is governed by privacy budget parameters that mathematically control the trade-off between privacy protection and analytical utility, as formalized in recent privacy computing research [93].
II-C3 Multi-Party Secure Computing (MPC)
MPC protocols enable multiple participants to jointly compute functions over their combined inputs while maintaining input privacy from each other. In distributed P2VAD scenarios, these cryptographic protocols allow different surveillance systems or organizations to collaboratively train anomaly detection models without sharing their respective video datasets. The underlying protocols ensure that each participant learns only the final computed result, such as model parameters or anomaly scores, without gaining access to other participants’ sensitive data. Liu et al. [94] identify MPC as a critical component for system-level privacy protection in VAD, particularly relevant for city-wide surveillance systems where different agencies must collaborate on anomaly detection while maintaining strict data sovereignty requirements.
II-C4 Federated Learning (FL)
FL extends the privacy preservation paradigm by enabling distributed model training without centralizing raw data. Rather than collecting video data from multiple sources into a central repository, federated protocols allow each participant to train local models on their private datasets and share only model parameters or gradients. The aggregation process combines these distributed updates into a global model that benefits from collective knowledge while preserving individual privacy. Doshi et al. [95] exemplify this approach through their FLVAD framework, which implements Transformer-based local models that process video data on trusted edge devices while participating in global model optimization through secure parameter sharing. Similarly, Al-Dujaili et al. [96] propose the CLAP framework, which employs common knowledge-based data segregation and local feedback to enable collaborative training without direct information exchange between clients, further enhancing privacy protection in federated VAD scenarios.
TABLE II: Technical Comparison of P2VAD methods with Non-Identifiable Elements.
Method Year Category Core Contributions Feature Type Processing Speed Privacy Mechanism
Kiryati et al. [71] 2008 Encrypted Videos Motion feature extraction from macroblock motion vectors Motion vectors Real-time Video compression
Biswas & Babu [97] 2014 Encrypted Videos Motion vector amplitudes in H.264/AVC compressed domain Motion vector amplitudes $>$150 fps H.264/AVC compression
Li et al. [69] 2014 Encrypted Videos Motion Intensity Count (MIC) feature using HEVC Motion vectors, coding units, PU patterns 1250 fps HEVC compression
Biswas & Babu [70] 2015 Encrypted Videos Enhanced feature set with motion vector orientation Motion vector amplitudes + orientation 90$\times$-250$\times$ speedup H.264/AVC compression
Sparse HOMV [98] 2015 Encrypted Videos Histogram of motion vectors with sparse representation HOMV sparse coefficients Not specified H.264 compression
Guo et al. [68] 2019 Encrypted Videos Parameter estimation in H.264 encrypted bitstreams Macroblock sizes, partition patterns, MV differences Not specified H.264 encryption
UBnormal [99] 2022 Synthesis Synthetic dataset generation for privacy preservation Synthetic video data Not specified Data synthesis
Schneider et al. [100] 2022 NVLCs Unsupervised anomaly detection using depth cameras Depth information Not specified Depth sensors
TeD-SPAD [101] 2023 Masking Object detection-based masking approach Masked visual features Not specified Object detection masking
Gaus et al. [81] 2023 NVLCs Region-based anomaly detection with infrared imaging Infrared features Not specified Infrared imaging
Yan et al. [102] 2024 Masking Advanced object detection-based masking Masked regions Not specified Object detection masking
@@ III P2VAD with Non-Identifiable Elements
Non-Identifiable Elements (NIE)-based P2VAD methods aim to prevent privacy and security risks by removing or abstracting identifiable information during data acquisition and preprocessing stages. The desensitized data is then used as input to the anomaly detection model. Based on the level of morphological abstraction, we classify existing NIE methods into three categories: encrypted videos, sensitive area masking/synthesis, and NVLCs videos, as shown in Fig. 3. In spite of privacy, such methods reduce data volume and prevent extraneous appearance information from interfering with model learning. However, these approaches have drawbacks, such as reliance on additional processing (e.g., compression, object detection) or the need for specialized acquisition equipment, which can significantly increase the cost of model deployment. The technical comparisons of NIE-based P2VAD methods are summarized in Table II.
Figure 3: Hierarchical taxonomy and related works of P2VAD with Non-Identifiable Elements (NIE). Note that the listed works include research papers, technical standards, and survey articles that embody related concepts and can inspire corresponding NIE-based P2VAD research directions.